Solve by the elimination method:
7r-9s=-35
9r+7s=85

Answer by mananth(16946)

7r-9s=-35
9r+7s=85
..
multiply first equation by -9 and equation2 by 7
-63r+81s=315
63r+49s=595
81s+49s=595+315
130s=910
s= 7
..
plug value of s in first equation
7r-9s=-35
7r-63=-35
7x=28
x= 4